OpenStudy (anonymous):

how do you find (y) from this equation 2x-y=-13 y+3x=-12

OpenStudy (anonymous):

First solve one of them for x, then plug in the expression you found into the other one for x and solve for y.

OpenStudy (anonymous):

first, add y to both sides to get 2x=y-13 then add 13 to both sides to get 2x+13=y

OpenStudy (anonymous):

for the second one do the same kind of thing, except just subtract 3x from both sides to get y=-12-3x

OpenStudy (anonymous):

2x-(-3x-12)=-13 2x+3x+12=-13 5x=-25 x=-5 y-15=-12 y=3

OpenStudy (anonymous):

I solved 2x-y=-13 but not exactly sure if i got it right? The (y) i got for it was y=13-2x

OpenStudy (dumbcow):

Hope i dont confuse things but i think elimination is much simpler for this problem: Add equations together 2x - y = -13 3x +y =-12 _____________ 5x + 0 = -25 divide by 5 on both sides x = -5 substitute back into original equation 2(-5) -y = -13 -10 - y = -13 Add 10 to both sides -y = -3 Flip signs y = 3

OpenStudy (anonymous):

If you are trying to find y, you first need to find x in terms of y from one equation, then plug it in. Or you can use elimination as cow suggests.

OpenStudy (anonymous):

Its asking to use the graph to solve the linear sytem. Then to check solution.

OpenStudy (dumbcow):

find the point where the 2 lines intersect, that is the solution check by substituting in the x,y values into one of the equations of the line

OpenStudy (anonymous):

okay now that makes more sense. thank you.